OpenStudy (lena772):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π (x - 7)2 (x - 3) What does the factor 1 over 3(x - 7) represent? The total surface area of the tank One third of the radius of the tank One third of the height of the tank The total area of the base of the tank

OpenStudy (anonymous):

question is missing something

OpenStudy (anonymous):

check if below line has typoes :- ``` What does the factor 1 over 3(x - 7) represent? ```

OpenStudy (lena772):

\[\frac{ 1 }{ 3 }(x-7)\]

Directrix (directrix):

For the volume, I got (1/3) *π * (x - 7)^ 2 * (x - 3)

OpenStudy (anonymous):

ohk... looking at the options, now it makes more sense

OpenStudy (anonymous):

(x-7) is ur radius right ?

OpenStudy (anonymous):

i should have asked, if \((x-3)\) is the height of tank, then \((x-7)\) is what ?

OpenStudy (anonymous):

height = x-3 radius = h-4 = x-3-4 = x - 7

OpenStudy (anonymous):

that means, (x-7) is the radius of tank so \(\frac{1}{3}(x-7)\) is One third the radius of tank

OpenStudy (lena772):

Ok

OpenStudy (lena772):

Thank you

OpenStudy (anonymous):

u wlc :)